OpenStudy (itsstribe):

Write an equation in point-slope form of the line that passes through given point and with the given slope m. (3,-4); m=6 Help, i'm so confused :/

OpenStudy (mathmale):

There are several equations one could use to describe a straight line. Since you're given one point on your line and the slope of your line, we'll use the "point-slope" form: \[y _{2}-y _{1}=m(x _{2}-x _{1})\]

OpenStudy (mathmale):

Let's go over that a bit more: You are given a point: (3,-4). This says that when x=3, y=-4. Please copy the "point-slope" equation above and replace \[y _{1}\] with -4 and replace \[x _{1}\] with 3.

OpenStudy (itsstribe):

Alright I will try..

OpenStudy (itsstribe):

2−-4=m(x2−3)

OpenStudy (mathmale):

\[y _{2}-y _{1}=m(x _{2}-x _{1})\] You are on the right track. It's just that those pesky subscripts 1 and 2 are a bit hard to type in.

OpenStudy (mathmale):

so I propose we go to the Draw utility: